OpenStudy (anonymous):

Write the equation in the standard form of an ellipse with center (h, k). 4x^2+3y^2+8x-6y-29=0

OpenStudy (anonymous):

4x^2 +8x + 4 + 3y^2 -6y +3 = 36 4(x+1)^2 + 3(y-1)^2 =36 (x+1)^2 / 9 + (y-1)^2 / 12 = 1

OpenStudy (anonymous):

got it?